In September 1993, The Times unilaterally lowered its price …

                 Price Average Daily Sale

            Pre- 09-93  Post-09-93   pre-09-93   post-09-93

Times       45  30  376836  448962

Guardian  45  45  420154  401705

Telegraph  45  45 1037375  1017326

Independent 50  50  362099  311046

2196464     2179039

Total daily newspaper sales (including a few other newspapers not mentioned in the slide) remained constant at about 2.5 million at both pre and post Sep-93.

Compute: Relevant cross-price elasticities of demand for Guardian, Telegraph, and Independent?

How do you ensure that the ceteris paribus assumption has been met here, although approximately?


1
Expert's answer
2021-07-29T11:24:02-0400

a)Cross price elasticity of demand: It is the measurement of the change in the quantity demanded of a commodity because of the change in the price of the related good. IT is measured using the below-mentioned formula:

"E_c =\frac{ ∆Q_A}{∆P_B}×\frac{P

_B}

{Q

_A}"


HereEc is the crossprice elasticity of demand

QA  is the change in quantity demanded for good A,

    QA  is the initial quantity demanded for good A,

PB is the change in price for good B

    PB is the initial price for good B


Cross-price elasticities of demand for Guardian;

The initial price of times = 45

Change in the price of times "= 45 - 30 = 15"

The initial quantity demanded of Guardian = 420154

Change in the quantity demanded of Guardian "= 420154 - 401705 = 18,449"

Putting in the value we get;

"E

_c

 = \frac{18449}

{15}



×\frac{45}

{420154}"

Ec = 0.13

Thus the cross-price elasticity of demand for Guardian is 0.13.


Cross-price elasticities of demand for Telegraph;

The initial price of times = 45

Change in the price of times = 45 - 30 = 15

The initial quantity demanded of Telegraph = 1037375

Change in the quantity demanded of Telegraph = 1037375 - 1017326 = 20,049

Putting in the value we get;

"E

_c

 =\frac{ 20049}

{15}



×{45}

{1037375}"

Ec = 0.058

Thus the cross-price elasticity of demand for telegraph is 0.058.

Cross-price elasticities of demand for Independent;

The initial price of times = 45

Change in the price of times = 45 - 30 = 15

The initial quantity demanded of Independent = 362099

Change in the quantity demanded of Independent = 362099 - 311046 = 51,053

Putting in the value we get;

"E

_c

 =\frac {51053}

{15}



×\frac{45}

{362099}"

Ec = 0.42

Thus the cross-price elasticity of demand for Independent is 0.42


b)Thus assumption ceteris paribus which means "other things remain constant" has been met here, because even after the constant price kept by Guardian, Telegraph, and Independent their quantity demanded fell because of the cross-price elasticities of demand. The Total daily newspaper sales (including a few other newspapers not mentioned in the slide) remained constant at about 2.5 million at both pre and post Sep-93 as a result of ceteris Paribus.